OpenStudy (anonymous):

Help me? :3 Verify the identity. cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x

OpenStudy (anonymous):

\[\cos 4x + \cos 2x= 2-2\sin^2 2x - 2\sin^2 x\]

OpenStudy (anonymous):

@jdoe0001 @whpalmer4 Okay, I know cos(2x)=1-2sin^2 x, so what does cos 4x equal?

OpenStudy (anonymous):

cos(2x + 2x) is one way to do it.

OpenStudy (anonymous):

Or cos[2(2x)]

OpenStudy (jdoe0001):

as already pointed out by @KinzaN \(\bf {\color{blue}{ cos(2\theta)=1-2sin^2(\theta)}}\\ \quad \\ \quad \\ cos(4x) + cos(2x) = 2 - 2 sin^2(2x) - 2 sin^2(x) \\ \quad \\ \implies cos[2({\color{green}{ 2x}})]+ cos(2x)\implies {\color{blue}{ 1-2sin^2({\color{green}{ 2x}})}}+{\color{blue}{ 1-2sin^2(x)}}\) simplify that, see what you get

OpenStudy (anonymous):

Oh. I was so close. Dadgummit.